identify the vertex focus axis of symmetry and directrix
x=-2y2-24y-76
Answer by lwsshak3(11628) 
 identify the vertex focus axis of symmetry and directrix
x=-2y2-24y-76
complete the square
x=-2(y^2+12y+36)+72-76
x=-2(y+6)^2-4
-2(y+6)^2=(x+4)
(y+6)^2=(-1/2)(x+4)
This is an equation of a parabola that opens leftward.
Its basic form: (y-k)^2=-4p(x-h)
For given equation:
vertex:(-4,-6)
axis of symmetry: y=-6
4p=1/2
p=1/8
directrix: x=31/8 (p-distance to the right of vertex on the axis of symmetry)
|
|
|
| |